### v4

# WellDunn: On the Robustness and Explainability of Language Models and Large Language Models in Identifying Wellness Dimensions

Abstract
Language Models (LMs) are being proposed for mental health applications where the heightened risk of adverse outcomes means predictive performance may not be a sufficient litmus test of a model’s utility in clinical practice. A model that can be trusted for practice should have a correspondence between explanation and clinical determination, yet no prior research has examined the attention fidelity of these models and their effect on ground truth explanations. We introduce an evaluation design that focuses on the robustness and explainability of LMs in identifying Wellness Dimensions (WDs). We focus on two existing mental health and well-being datasets: (a) Multi-label Classification-based MultiWD, and (b) WellXplain for evaluating attention mechanism veracity against expert-labeled explanations. The labels are based on Halbert Dunn’s theory of wellness, which gives grounding to our evaluation. We reveal four surprising results about LMs/LLMs: (1) Despite their human-like capabilities, GPT-3.5/4 lag behind RoBERTa, and MedAlpaca, a fine-tuned LLM on wellxplain fails to deliver any remarkable improvements in performance or explanations. (2) Re-examining LMs’ predictions based on a confidence-oriented loss function reveals a significant performance drop. (3) Across all LMs/LLMs, the alignment between attention and explanations remains low, with LLMs scoring a dismal 0.0. (4) Most mental health-specific LMs/LLMs overlook domain-specific knowledge and undervalue explanations, causing these discrepancies. This study highlights the need for further research into their consistency and explanations in mental health and well-being.
WellDunn: On the Robustness and Explainability of Language Models and Large Language Models in Identifying Wellness Dimensions
Seyedali Mohammadi1*, Edward Raff1,2, Jinendra Malekar3, Vedant Palit4, Francis Ferraro1, Manas Gaur1 1UMBC, MD, USA 2Booz Allen Hamilton 3USC, SC, USA 4IIT, Kharagpur, India *mohammadi@umbc.edu
1 Introduction
According to the National Institute of Mental Health (nih2023MentalIllness), over 20% of US adults have experienced mental illnesses, prompting the government to allocate \$280 billion to address unmet mental health service needs (whitehouseReducingEconomic). This highlighted the need to leverage AI (particularly LMs/LLMs) for mental health, as they can potentially decrease costs and increase the accessibility of mental health services. However, vigilance is crucial regarding the potential risk of LMs/LLMs arising from low-confidence predictions and correct predictions with wrong explanations.
Figure 1: Motivating Example from WellXplain dataset. Expert annotators categorize user posts into four WD classes and justify their choice by highlighting pertinent parts of the text. In LM or LLM classification tasks, the goal is to identify one of the labels (1: Physical, 2: Intellectual and Vocational, 3: Social, 4: Spiritual and Emotional) based solely on relevant cues in the post. The cues are the explanations.
Motivated by this longer-term goal of safe deployment of NLP-based mental health systems, we propose evaluation schemes examining the consistency in LM’s attention (and LLM’s attention where the attention is accessible) with ground-truth explanations [^1] In this context, we are using ’explanation’ that refer to ’text-span explanations’ which are tokens/spans of text that are relevant for determining class labels. and confidence in predictions. Our insight is that a model’s attention in disagreement with physician assessment is unlikely to be accepted, regardless of predictive accuracy. Indeed, such a scenario implies the model has learned some shortcut or correlative signal instead.
Figure 2: WellDunn workflow: MultiWD task (L) and WellXplain task (R). The architecture includes shared steps: (1) Fine-tuning of general purpose and domain-specific LMs for extracting data representations, followed by (2) feeding them into a feed-forward neural network classifier (FFNN). Two loss functions assess LMs’ robustness: Sigmoid Cross-Entropy(SCE) and Gambler’s Loss(GL). Singular Value Decomposition (SVD) and Attention-Overlap (AO) Score assess the explainability. In: Input, and Out: Output. WellDunn Benchmarking Box: This middle rectangle highlights the components of the benchmark system, which includes steps of (1) Fine-tuning and (2) FFNN classifier, as well as Robustness and Explainability components. The Left and right dotted rectangles grouped the components for the MultiWD and WellXplain tasks, respectively. In the case of Task 1, the input (text post) is fed into the MultiWD task, and the model produces an output (prediction) in terms of various WDs like PA, IA, SA, etc. For Task 2, the input (text post) is also fed into the WellXplain task, which produces output (prediction) along with corresponding explanations. Note that in the instruction (training), we provide both input and output, but in the evaluation (test), we provide the input.
We present an evaluation framework, acronymized as WellDunn, which examines 11 LMs/LLMs using two domain-grounded datasets annotated with the causes of deteriorating wellness in individuals. These datasets are important because they consist of user-generated content from individuals expressing signs of depression, bipolar disorder, anxiety, suicide, schizophrenia, or comorbidities caused by the decline in their wellness. The MultiWD222 https://github.com/drmuskangarg/MultiWD dataset (sathvik2023multiwd) assigns six interconnected Wellness Dimensions (WDs)—Physical, Intellectual, Vocational, Social, Spiritual, and Emotional—to each textual post (crawled from Reddit’s posts) based on Halbert L. Dunn’s classification (dunn1959high; sathvik2023multiwd). This dataset frames the task as a multi-label classification, evaluating LMs/LLMs predictive performance in contexts where WDs are interdependent  (hallerod2013multi). The WellXplain [^3] https://github.com/drmuskangarg/WellnessDimensions/ dataset (garg2024wellxplain; liyanage2023augmenting) assigns a single WD to each textual post, with annotations explaining the reasons behind the label. Figure 1 presents an example from WellXplain, where an LM/LLM predicts a WD and offers an explanation, highlighting the text that captures the model’s attention.
Welldunn Evaluation Criteria:
We utilize traditional evaluation metrics along with supplementary ones, including SVD rank, Attention-Overlap score, and Attention Maps. The SVD rank assesses the focus of attention in LMs [^4] LLMs’ internal machinery is not as transparent as LMs’. while the Attention-Overlap score measures the extent to which the model’s attention aligns with ground truth explanations in the WellXplain dataset. Figure 2 illustrates the procedure of WellDunn.
Findings:
Our empirical research into LMs and LLMs for mental health and well-being revealed several key findings: (a) domain-specific LMs/LLMs performed within 1% of general-purpose models; on average, general-purpose LMs showed a 1.3% improvement in performance over domain-specific LMs/LLMs. (b) general-purpose LMs exhibited higher confidence in their predictions compared to domain-specific models. After retraining four general-purpose and three domain-specific LMs with a confidence-oriented loss function—gambler’s loss (a variant of sigmoid cross-entropy)— general-purpose LMs exhibited 6.3% higher confidence and significantly better attention compared to domain-specific LMs. The decrease in scores is attributed to LMs abstaining from making low-confidence predictions. (c) general-purpose LMs demonstrated more focused attention than domain-specific LMs, including LLAMA and MedAlpaca. In an inter-model comparison on WellXplain, LLMs underperformed by 32.5% in MCC compared to vanilla RoBERTa, which also demonstrated higher confidence.
Takeaway: These findings challenge assumptions about the efficacy of larger models and the value of fine-tuning in mental health applications. These gaps lead to incorrect and misleading explanations when these models are queried for causes of mental health issues, undermining their reliability and clinical utility. The attention overlap score of 0.0 for LLAMA and MedAlpaca, along with the significant gap between SVD rank and the average length of explanations, supports our inferences and demonstrates significant failures can still occur.
Note: Attention as a medium of explanation is debatable, as inferred from prior works by bibal2022attention; jain2019attention and wiegreffe2019attention. However, in these studies, the datasets did not have explicit expert-provided explanations, which can be used to cross-check the overlap between high-attention words and natural language explanations. As in this research, we have a dataset with natural language explanations; we consider attention a medium of explanation.
2 Related Work
AI in Mental Health: Previous studies in the convergence of AI and mental health concentrated on creating or improving machine learning and deep learning algorithms to identify mental health conditions (MHCs) or assess their severity (lin2020sensemood; cao2020building; lin2017detecting; haque2021deep). However, minimal attention has been dedicated to ensuring these AI-driven models’ robustness and explanatory capabilities. As a result, researchers and practitioners lack insight into whether these models emphasize the correct clinically relevant terms to make decisions and whether they are made with confidence.
To overcome this challenge, efforts have been made to create knowledge-grounded, expert-curated datasets incorporating clinical expertise. These datasets utilize clinical knowledge in various forms, such as human experts acting as crowd workers, e.g. shen2017depression, CLPsych by coppersmith2015clpsych, mental health lexicons (gaur2019knowledge), and clinical practice guidelines (gupta2022learning; zirikly2022explaining). A recent study by garg2023mental has enumerated 17 classification datasets focused on mental health outcomes, including suicide risk, depression, mental health, stress, and emotion. Various domain-specific and general-purpose LMs have been trained on these datasets. However, the robustness and attention of these models have not been thoroughly examined. This study addresses this gap by adapting WellXplain’s clinically validated explanations for comparative analysis alongside attention mechanisms so that we can test model attention’s alignment with a causal determinant.
Wellness Dimensions: The severity of MHCs and their cormorbities varies among individuals (coppersmith2021individual). Despite knowledge-grounded datasets, LMs face challenges in generalizing effectively (harrigian2020models). This difficulty arises from overlooking signs of mental disturbances that can trigger sub-clinical depression and progress to clinical depression over extended periods if left undetected. These signs go beyond the traditional psycholinguistic assessment of natural language, which involves using lexicons like LabMT (reagan2018labmtsimple), ANEW (bradley1999affective), and LIWC (pennebaker2001linguistic). There’s a rising interest in using WDs to advance mental health research with LMs (liyanage2023augmenting). This study is the first to use LMs in mental health, focusing on the model’s attention and confidence in predicting WDs.
3 Datasets
Dataset #Sample Avg. words/post MultiWD 3281 632 WellXplain 3092 112
Table 1: Basic statistics of MultiWD and WellXplain datasets: #Sample and Avg. words/post represent the number of samples (each sample includes a post and its six labels) and the average number of words per post respectively.
We utilize two expert-annotated and domain-grounded datasets, MultiWD (sathvik2023multiwd) and WellXplain (garg2024wellxplain), which are based on Halbert Dunn’s seminal wellness concepts (dunn1959high). To the best of our knowledge, MultiWD and WellXplain are the only datasets available for WDs. Task 1 (MultiWD) involves multi-label classification, while Task 2 (WellXplain) involves multi-class classification with expert annotator explanations, as summarized in Table 1. These datasets encompass six dimensions of wellness: Physical Aspect (PA), Intellectual Aspect (IA), Vocational Aspect (VA), Social Aspect (SA), Spiritual Aspect (SpA), and Emotional Aspect (EA). The definitions for these aspects by sathvik2023multiwd can be found in LABEL:WDsDefinitions.
Task 1:
The MultiWD dataset consists of 3281 instances, each comprising a text post and six distinct binary labels indicating whether a particular WD is present (1) or absent (0). The posts are crawled from Reddit’s two most prominent mental health forums: r/Depression and r/SuicideWatch (sathvik2023multiwd).  LABEL:table:Conditions (LABEL:Appendix_D) presents the number of posts where a user explicitly refers to a mental health condition and specifies one or more wellness aspects impacted. In  LABEL:table:Conditions (LABEL:Appendix_D), the users primarily mention depression, anxiety, and suicide as prominent MHCs impacting their social and emotional wellness. This corresponds to our intention of utilizing WD as a preliminary task to fine-tune LMs before engaging in binary mental health classification (refer to Figure 1).
Figure 3: Merging of WDs in MultiWD.The expert annotators suggest merging WDs based on their experience and literature (bart2018assessment).
Subtask of Merging WDs in MultiWD: The WDs include many overlapping wellness tenets and individual proponents, making them unique. It is important to exercise the performance of LMs by merging WDs. The expert annotators suggest merging WDs based on their experience and literature (bart2018assessment). For instance, spiritual wellness is closely related to emotional wellness (in this specific clinical framework). Thus, we merge the most related classes in MultiWD to explore how performance changes in an easier (4 classes) vs harder (5 to 6 classes) setting. Figure 3 shows the merging of WDs.
Task 2:
The WellXplain dataset comprises 3092 instances from r/Depression and r/SuicideWatch. Each instance includes a text post, accompanying explanatory information, and a specific WD aspect assignment. LABEL:table:Conditions (LABEL:Appendix_D) shows the presence of depression and anxiety as the top two MHCs expressed in the dataset impacting spiritual, emotional, social, and physical wellness. “Explanations” in WellXplain refer to the textual cues considered by annotators when determining the classification into one of the four predefined categories: (1) PA, (2) IVA, (3) SA, and (4) SpEA.
4 WellDunn: Methodology and Evaluation
Domain-specific and General-purpose LMs: We consider two distinct categories of models for the task of WD identification – general models and domain-specific models. The general models under consideration include BERT (devlin2018BERT), RoBERTa (liu2019roBERTa), Xlnet (yang2019xlnet), and ERNIE (sun2019ernie). Additionally, we incorporate domain-specific models, namely ClinicalBERT (alsentzer2019publicly), MentalBERT (ji2021mentalBERT), and PsychBERT (vajre2021psychbert), to further explore their applicability within the mental health domain.